OpenStudy (anonymous):

A pair of sunglasses costs the retailer $80 per pair. At what price should the retailer mark them so he can sell them at a 20% discount off the selling price and still make a 20% profit on his cost?

OpenStudy (anonymous):

cost price of glasses = 80 he wants to make 20% profit ie. 20% of 80 = 0.20*80 = 16 selling price = 80+16 = 96 let the marked price = x discount = 20% of x = 0.20x selling price = marked price - discount = x - 0.20x = 0.80x ATQ 0.80x = 96 x = 96/0.80 x = 120 so the marked price will be $120